# Calibrating Verbalized Confidence with Self-Generated Distractors

### v1

@@ 3.2 Results
Table 1: TriviaQA results. We evaluate Expected Calibration Error (ECE), Brier score (BS), and area under the ROC curve (AUC). In each column, we bold the best result and underline results not significantly worse under a paired test ($\alpha=0.05$; see Appendix B.6 for tests).
@@
Short-form QA.  Tables 1 and 2 report results for TriviaQA and SimpleQA. On TriviaQA, DiNCo outperforms the best baseline, SC, by an average ECE of 0.099. On SimpleQA, DiNCo outperforms the best baseline, MSP, by an average ECE of 0.092. While the better baseline varies by dataset, DiNCo consistently achieves the strongest calibration. MSP is sometimes a competitive baseline (e.g. AUC 0.800 surpasses DiNCo at 0.786 on SimpleQA with GPT-4.1), but this often does not hold across metrics (e.g. MSP has an ECE of 0.263 in the same setting, heavily underperforming DiNCo at 0.089) or across settings (e.g. MSP underperforms DiNCo on AUC in TriviaQA by an average of 0.062). Moreover, we note that the effectiveness of MSP relies on answers having a canonical form, preventing its generalization to long-form settings (Farquhar et al., 2024). We highlight that NVC outperforms SC (e.g. by an ECE of 0.049 and 0.060 on TriviaQA and SimpleQA, respectively) despite only leveraging coherence in validation and not in generation (Section 2.3). Nonetheless, DiNCo is more consistently calibrated than NVC (in particular on AUC, e.g. 0.786 DiNCo vs. 0.729 NVC with GPT-4.1 on SimpleQA), empirically supporting our motivation in Section 2.3 for integrating coherence in generation (SC) and validation (NVC) into DiNCo. In the black-box setting on SimpleQA, DiNCo continues to do well (e.g. outperforming the baselines on ECE), but it tends to fall behind DiNCo with logit access, underscoring the benefit of leveraging token probabilities for calibration.
@@ 4 Discussion and Analysis
Figure 4: Scaling self-consistency does not close the gap with DiNCo.
@@
While our main experiments in Section 3 were controlled for the inference budget of each method, here we examine whether simply scaling the inference budget of self-consistency allows it to recover the calibration of DiNCo. Fig. 4 shows that self-consistency alone is unable to reach the performance of DiNCo (using 5 distractors and 5 self-consistency samples as in Section 3) even when scaling up to 100 samples, demonstrating that the effectiveness of DiNCo comes from leveraging coherence in both generation and validation, which is not matched by scaling the generation axis alone.
@@
Our main experiments in Tables 1, 2 and 3 showed comparisons with SC and NVC, which ablate NVC and SC, respectively, from DiNCo. Here, to understand how necessary access to an NLI model is, we ablate the NLI component, which is used to downweight distractors that overlap with the main claim or other distractors (Section 2.3, Fig. 2). In Table 5, we see that performance substantially decreases without NLI-based weighting, emphasizing the utility of an off-the-shelf NLI model to account for claim redundancy.
@@ 5 Related Work
Considering Multiple Answers for Verbalized Confidence.  The approach most related to DiNCo is to have the model consider several candidates within a single prompt and assign confidences to them (Tian et al., 2023; Kadavath et al., 2022; Zhang et al., 2024b). Similarly, Chhikara (2025) samples answers and presents them in the context as distractors. A subtle but crucial distinction between these methods and DiNCo is that if we present all the candidates together, we become unable to gauge the probabilistic coherence of the confidence estimates, i.e. whether they form a valid probability distribution, since LLMs can satisfy probabilistic coherence via simple arithmetic. In Section 2.2, we show that the probabilistic coherence of confidence estimates is correlated with answer correctness. Instruction-tuned LLMs have a tendency to assert confidence even when it is undue (OpenAI, 2023; Leng et al., 2025; Sun et al., 2025; Xu et al., 2025), leading joint prompting to suffer similar issues of overconfidence as vanilla prompting. With independent prompting, we can expose and account for inconsistencies in self-declared knowledge. In Section 3, we empirically verify that our method leads to better calibration than joint prompting for verbalized confidence.
@@ B.3 LLM-as-a-Judge
We score model responses for short-form QA (TriviaQA and SimpleQA) using an LLM judge rather than lexical matching for robust evaluation (Wei et al., 2024). Our biography generation task is evaluated with FActScore (Min et al., 2023), which uses a strong LLM for atomic claim decomposition and verification; we use GPT-4.1. We use Llama-3.1-8B-Instruct (Dubey et al., 2024) as the judge on TriviaQA, and GPT-4.1 as the judge on SimpleQA. On a sample of 100 questions from each dataset, we compared LLM judgments with human judgments. We used responses from Qwen3-8B on TriviaQA, and responses from GPT-4.1 on SimpleQA. The rate of agreement was 96/100 and 99/100, respectively. Upon reviewing disagreements, we found that in most cases there was genuine room for interpretation. For example, for the question “In The Living Daylights what did Carla keep in her cello case?”, the target answer is “A machine gun” while the model answer was “A gun”. It is unclear whether the model answer has the desired specificity. As another example, for the question “In Charles Dickens’ “Great Expectations”, who or what was Abel Magwitch?”, the target answer is “Convict” while the model answer was “A convict and the main benefactor of Pip”. The model answer contains the answer but contains more information, and without external information, it is impossible to determine whether the model’s answer is correct. Overall, given the high agreement on unambiguously gradable questions, we deem it safe to adopt LLM-as-a-judge as a reliable evaluator in our experiments.
@@ B.4 FActScore Example
B.5 NLI Model
We use the NLI model DeBERTa-v3-base-mnli-fever-anli (He et al., 2021). [^8] https://huggingface.co/MoritzLaurer/DeBERTa-v3-base-mnli-fever-anli
B.6 Significance Testing
@@ Appendix C Methods
C.1 Pseudo-Beam Search
@@ C.2 Self-consistency
Appendix D Results
D.1 DiNCo with Verbalized Numerical Confidence
@@
In Tables 7 and 8, we expand on Tables 1 and 3 and report results on TriviaQA and biography generation when replacing $\rm P(True)$} with verbalized numerical confidence (see Appendix B.2 for prompts). For SimpleQA, the black-box variants of our methods in Table 2 use verbalized numerical confidence rather than $\rm P(True)$}.

### v2

@@ 3.2 Results
Table 1: TriviaQA results. We evaluate Expected Calibration Error (ECE), Brier score (BS), and area under the ROC curve (AUC). In each column, we bold the best result and underline results not significantly worse under a paired test ($\alpha=0.05$; see Appendix B.5 for tests). For readability of this table, we leave Qwen3-32B results to Appendix D.5, where we verify that the effectiveness of DiNCo extends to larger scales of open-source models.
@@
Short-form QA.  Tables 1 and 2 report that on TriviaQA and SimpleQA, DiNCo outperforms the best baseline, MSP, by an average ECE of 0.077 and 0.092, respectively. While MSP is a competitive baseline (e.g. AUC 0.800 surpasses DiNCo at 0.786 on SimpleQA with GPT-4.1), this often does not hold across metrics (e.g. MSP has an ECE of 0.263 in the same setting, heavily underperforming DiNCo at 0.089) or across settings (e.g. MSP underperforms DiNCo on AUC in TriviaQA by an average of 0.062). Most importantly, the effectiveness of MSP relies on answers having a canonical form, restricting its usage to multiple-choice or short-form questions and preventing its generalization to long-form settings (Farquhar et al., 2024). We highlight that NVC outperforms SC (e.g. by an ECE of 0.049 and 0.060 on TriviaQA and SimpleQA, respectively) despite only leveraging coherence in validation and not in generation (Section 2.3). Nonetheless, DiNCo is more consistently calibrated than NVC (in particular on AUC, e.g. 0.786 DiNCo vs. 0.729 NVC with GPT-4.1 on SimpleQA), empirically supporting our motivation in Section 2.3 for integrating coherence in generation (SC) and validation (NVC) into DiNCo. In the black-box setting on SimpleQA, DiNCo continues to do well (e.g. outperforming the baselines on ECE), but it tends to fall behind DiNCo with logit access, underscoring the benefit of leveraging token probabilities for calibration. In Appendix D.6, we evaluate Qwen3-32B on BioASQ (Krithara et al., 2023), extending these findings to the biomedical domain where expert knowledge is required.
@@ 4 Discussion and Analysis
[…]
Figure 4: Scaling self-consistency does not close the gap with DiNCo. FLOP counts are with Qwen3-8B on 1000 TriviaQA questions. Since each distractor requires a verbalized confidence step, DiNCo costs 32% more than SC. However, due to diminishing returns, scaling SC up to even 100 samples negligibly improves performance (despite costing 7.6 times as much as DiNCo), justifying the slightly higher cost of DiNCo to achieve stronger calibration. We remark that the cost of the lightweight NLI model (184M parameters) used in DiNCo and SC is negligible, making up less than 1% of the total FLOP count.
@@
While our main experiments in Section 3 were approximately controlled for the inference budget of each method, here we examine whether simply scaling the inference budget of self-consistency allows it to recover the calibration of DiNCo. Fig. 4 shows that self-consistency alone is unable to reach the performance of DiNCo (using 5 distractors and 5 self-consistency samples as in Section 3) even when scaling up to 100 samples. Although DiNCo uses 32% more FLOPs than SC in the context of Section 3, scaling SC up to match and far exceed this slightly higher FLOP count fails to further improve calibration, demonstrating that the effectiveness of DiNCo comes from leveraging coherence in both generation and validation, which is not matched by scaling the generation axis alone.
In Appendix D.7, we show that combining distractor generation and validation into one step maintains similar calibration to the original two-step version while reducing cost to only 10% more than SC. In Appendix C.1, we motivate the feasibility of generating distractors using a smaller model to reduce cost below even that of SC.
@@
Our main experiments in Tables 1, 2 and 3 showed comparisons with SC and NVC, which ablate NVC and SC, respectively, from DiNCo. Here, to understand how necessary access to an NLI model is, we ablate the NLI component, which is used to downweight distractors that overlap with the main claim or other distractors (Section 2.3, Fig. 2). In Table 5, we see that performance substantially decreases without NLI-based weighting, emphasizing the utility of an off-the-shelf NLI model to account for claim redundancy. In Appendix D.3, we show that DiNCo is robust to the specific choice of the NLI model.
@@ 5 Related Work
Considering Multiple Answers for Verbalized Confidence.  The approach most related to DiNCo is to have the model consider several candidates within a single prompt and assign confidences to them (Tian et al., 2023; Kadavath et al., 2022; Zhang et al., 2024b; Chhikara, 2025). A subtle but crucial distinction between these methods and DiNCo is that if we present all the candidates together, we become unable to gauge the probabilistic coherence of the confidence estimates, i.e. whether they form a valid probability distribution, since LLMs can satisfy probabilistic coherence via simple arithmetic. In Section 2.2, we show that the probabilistic coherence of confidence estimates is correlated with answer correctness. Instruction-tuned LLMs have a tendency to assert confidence even when it is undue (OpenAI, 2023; Leng et al., 2025; Sun et al., 2025; Xu et al., 2025), leading joint prompting to suffer similar issues of overconfidence as vanilla prompting. With independent prompting, we can expose and account for inconsistencies in self-declared knowledge. In Section 3, we empirically verify that our method leads to better calibration than joint prompting for verbalized confidence.
@@ 6 Conclusion
Acknowledgments
The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper.
@@ B.3 LLM-as-a-Judge
We score model responses for short-form QA (TriviaQA and SimpleQA) using an LLM judge rather than lexical matching for robust evaluation (Wei et al., 2024). Our biography generation task is evaluated with FactScore (Min et al., 2023), which uses a strong LLM for atomic claim decomposition and verification; we use GPT-4.1. We use Llama-3.1-8B-Instruct (Dubey et al., 2024) as the judge on TriviaQA, and GPT-4.1 as the judge on SimpleQA. On a sample of 100 questions from each dataset, we compared LLM judgments with human judgments from one author of this paper. We used responses from Qwen3-8B on TriviaQA, and responses from GPT-4.1 on SimpleQA. The rate of agreement was 96/100 and 99/100, respectively. Upon reviewing disagreements, we found that in most cases there was genuine room for interpretation. For example, for the question “In The Living Daylights what did Carla keep in her cello case?”, the target answer is “A machine gun” while the model answer was “A gun”. It is unclear whether the model answer has the desired specificity. As another example, for the question “In Charles Dickens’ “Great Expectations”, who or what was Abel Magwitch?”, the target answer is “Convict” while the model answer was “A convict and the main benefactor of Pip”. The model answer contains the answer but contains more information, and without external information, it is impossible to determine whether the model’s answer is correct. Overall, given the high agreement on unambiguously gradable questions, we deem it safe to adopt LLM-as-a-judge as a reliable evaluator in our experiments.
@@ B.4 FactScore Example
B.5 Significance Testing
@@ Appendix C Methods
C.1 Examples, Analysis, and Discussion of Distractors
Examples.
Building on the example in Fig. 2, we provide more examples of distractors generated by LLMs using different methods of self-generating distractors: beam search on a vanilla QA prompt (Table 7), beam search on a prompt eliciting one distractor for a given claim (Table 8), pseudo-beam search on a vanilla QA prompt (Table 9(a)), and a prompt eliciting a list of candidate answers (Table 9(b)); see Sections 2.3 and 3 for further descriptions and Appendix B.2 for prompts. For each distractor, we also report the verbalized confidence, $w_{\rm unique}$, and $w_{\rm contra}$, which underlie the computation of $\beta(C)$ in Eq. 4 and measure plausibility, uniqueness, and counterfactuality, respectively. We report these quantities under the experimental settings in Section 3.
Quantitative Analysis.
We analyze these attributes of self-generated distractors over a full dataset. Fig. 5 shows that beam search on a vanilla QA prompt leads to distractors with high mean values of $\rm P(True)$} (indicating plausibility), $w_{\rm unique}$ (indicating low repetitiveness), and $w_{\rm contra}$ (indicating opposition to the main claim). In other words, beam search on a vanilla QA prompt enables efficient and diverse exploration of alternative claims, despite not explicitly prompting for distractors. We summarize this analysis for the other distractor generation methods in Fig. 6, where we consistently observe high average values of verbalized confidence, $w_{\rm unique}$, and $w_{\rm contra}$, suggesting that many self-generated distractors fulfill the desiderata of being plausible, unique, and counterfactual.
Discussion.
Identifying entailment and contradiction relationships may be a nontrivial task in general, such as in cases requiring domain expertise. DiNCo does not dictate the choice of the NLI model, so in such cases, it may be worth using a stronger NLI model or even the generator LLM itself. While this reliance on the NLI capabilities of the same LLM we are attempting to calibrate may seem circular, we argue that NLI is a tractable and modular task that is largely solved, in contrast to LLM calibration. In this paper, we empirically show that a lightweight off-the-shelf NLI model suffices to achieve strong calibration across short-form and long-form generation settings. In Appendix D.6, we show that the same lightweight NLI model continues to suffice in the biomedical domain despite its highly technical nature. In Appendix D.3, we show that DiNCo is robust to the choice of the NLI model.
DiNCo uses the generator LLM to generate its own distractors to relieve the need for external models. However, one concern is that a model’s epistemic uncertainty may limit its ability to generate plausible, high-quality distractors. Fortunately, our quantitative analysis (Figs. 5 and 6) provides empirical evidence of the plausibility of generated distractors. To offer an intuitive explanation of this behavior, we note that the “plausibility” of a claim depends on the model assessing this plausibility. If a model is epistemically uncertain, the distractors it generates may be implausible to an expert model that is epistemically certain, but this is acceptable because we are interested in whether the distractors are plausible to the epistemically uncertain model so that we can leverage the phenomenon of suggestibility under epistemic uncertainty (Section 2.1). As an extreme case of this, it is even acceptable for a distractor to be factual, as long as it is counterfactual to the main claim. As shown by the examples in Table 9(b), we want to know whether the model can distinguish correct and incorrect claims, regardless of whether the distractors happen to be correct.
The discussion above for the well-suitedness of a model to generate its own distractors also hints at the potential for distractors to be generated by a smaller model to reduce computational cost. Although we do not investigate this idea in this paper, we expect the task of generating somewhat plausible distractors to be much easier than knowing the correct answer. Considering the example in Table 9, the information required to correctly answer the question is obscure, but anyone who understands the question can produce distractors that would likely be indistinguishable from the correct answer to someone else who does not know the answer to the question. This suggests that a model with a smaller capacity than the model we seek to calibrate could be used for efficient distractor generation with little degradation in distractor quality. Outside of calibration, other methods such as contrastive decoding (Li et al., 2023) leverage a similar intuition that weaker models can produce plausible but incorrect candidate generations.
Table 7: Example distractors and their attributes. Qwen3-8B on TriviaQA. Main answer and distractors generated using beam search on a vanilla QA prompt (Appendix B.2.1).
Question
In The Living Daylights what did Carla keep in her cello case
True answer
A machine gun
Main answer VC $\beta(C)$ NVC
A gun 1.00 5.91 0.17
Distractor VC $w_{\rm unique}$ $w_{\rm contra}$
A bomb 1.00 1.00 1.00
A jar of honey 0.12 1.00 1.00
A set of keys 0.92 1.00 1.00
A jar of spiders 0.82 0.99 1.00
A stash of drugs 1.00 1.00 1.00
A pair of binoculars 0.01 1.00 1.00
A jar of pickles 0.99 0.50 1.00
A jar of pickled gherkins 1.00 0.51 1.00
A secret weapon 0.97 0.54 0.10
Table 8: Example distractors and their attributes. Qwen3-8B on FactScore (biography generation). The main claim is extracted from the generated biography, and the distractors are generated using beam search on a prompt eliciting one distractor given the main claim (Appendix B.2.2).
Main claim (entity: Kalki Koechlin) VC $\beta(C)$ NVC
She played Nandini in the film Chandni Chowk to China. 0.49 1.48 0.33
Distractor VC $w_{\rm unique}$ $w_{\rm contra}$
She played Chandni in the film Chandni Chowk to China. 0.07 0.50 0.98
She played Nandini in the film Chandani Chowk to China. 0.81 0.98 0.00
She played Nandini in the film Chandni Chowk to Russia. 0.45 0.99 1.00
She played Nandini in the film Titanic. 0.00 1.00 1.00
*** She played Chandni in the film *Chandni Chowk to China*. 1.00 0.50 0.98
She played Nandini in the film Chandni Chowk to Italy. 0.01 0.99 1.00
She played Nandini in the film Dangal. 0.00 1.00 0.99
She played Nandini in the film Chandni Chowk to Canada. 0.00 0.99 1.00
Table 9: Example distractors and their attributes. GPT-4.1 on SimpleQA. Table 9(a) shows examples using pseudo-beam search on a vanilla QA prompt to generate distractors. Table 9(b) shows examples in the black-box setting using a single generation to a prompt asking for a list of candidate answers, and the verbalized confidence is verbalized numerical confidence rather than $\rm P(True)$} (we still elicit confidence for each candidate answer independently). See Appendix B.2.1 for prompts.
Question
For how many years did Mohamed Abdelaziz Djaït serve as the Mufti of the Republic of Tunisia?
True answer
3
((a)) Pseudo-beam search.
Main answer VC $\beta(C)$ NVC
17 years 0.85 6.19 0.14
Distractor VC $w_{\rm unique}$ $w_{\rm contra}$
Twenty years 0.78 0.52 0.99
Two years 0.62 1.00 1.00
12 years 0.22 1.00 1.00
22 years 0.12 0.54 1.00
Three years 0.90 1.00 1.00
11 years 0.90 1.00 1.00
Six years 0.73 1.00 1.00
Eight years 0.78 1.00 1.00
Four years 0.73 1.00 1.00
((b)) Black-box prompting.
Main answer VC $\beta(C)$ NVC
2 1.00 9.56 0.10
Distractor VC $w_{\rm unique}$ $w_{\rm contra}$
3 1.00 1.00 0.99
1 1.00 1.00 0.99
4 1.00 1.00 1.00
5 0.80 1.00 1.00
6 1.00 1.00 1.00
7 1.00 1.00 1.00
8 1.00 1.00 1.00
9 1.00 1.00 1.00
10 0.80 1.00 1.00
Figure 5: Pairwise correlations (Pearson’s $r$, Spearman’s $\rho$) between distractor attributes: $\rm P(True)$}, $w_{\rm unique}$, and $w_{\rm contra}$. Due to the large numbers of points, we depict joint distributions by plotting the density using kernel density estimation, with darker shades indicating more points. The diagonals plot the marginal distributions. These results are with Qwen3-8B on TriviaQA, using beam search on a vanilla QA prompt (Appendix B.2.1) to generate distractors. In other words, we do not explicitly prompt the LLM to generate incorrect answers, motivating our investigation of the plausibility, uniqueness, and counterfactuality of the distractors. We find that $w_{\rm unique}$ and $w_{\rm contra}$ are positively correlated, indicating that distractors that contradict the main claim tend to be repeated fewer times within the distractor set, suggesting a diverse exploration of the set of alternative claims. On the other hand, $w_{\rm unique}$ and $w_{\rm contra}$ are negatively correlated with $\rm P(True)$}. Interpreting $\rm P(True)$} as a measure of plausibility, this means that plausible claims are more repeatedly generated and are more in agreement with the main claim. Nevertheless, these negative correlations are weak, and the fact that the mean values of $\rm P(True)$} and $w_{\rm contra}$ are both high indicates that a substantial number of plausible distractors are being generated.
Figure 6: Statistics of distractor attributes (verbalized confidence, $w_{\rm unique}$, and $w_{\rm contra}$) for different methods of self-generating distractors. We report their means and Spearman correlations. Although the second method (beam search on a prompt eliciting one distractor for a given claim) has no correlation between $w_{\rm unique}$ and $w_{\rm contra}$ unlike the other methods, we find that all methods have high average values of verbalized confidence, $w_{\rm unique}$, and $w_{\rm contra}$, suggesting that many self-generated distractors fulfill the desiderata of being plausible, unique, and counterfactual. Notably, the fourth method (prompting for a list of candidate answers) has the highest $w_{\rm unique}$ and $w_{\rm contra}$, showing the potential benefit of conditioning on previously generated distractors rather than generating them independently. We note that the fourth method uses verbalized numerical confidence rather than $\rm P(True)$} due to being in the black-box setting (we still elicit confidence for each candidate answer independently).
C.2 Pseudo-Beam Search
@@ C.3 Self-consistency
Appendix D Ablations and Results
D.1 Preliminary Study Extended
To extend the preliminary study presented in Section 2.2 comparing confidence levels on correctly and incorrectly answered questions, we provide the same analysis for the other two datasets in our main experiments (Section 3). SimpleQA (Wei et al., 2024) is a short-answer QA dataset similar to TriviaQA except designed to challenge frontier models. We accordingly evaluate GPT-4.1 on SimpleQA, thus extending our result to the scale of frontier models. Biography generation is evaluated with FactScore (Min et al., 2023). Similar to Fig. 3, Fig. 7 finds higher total confidence (over the main claim and distractors) on incorrectly answered questions, consistent with our hypothesis that LLMs are more susceptible to suggestibility when epistemically uncertain.
Figure 7: Distributions of total confidence over correct and incorrect answers, showing the same trends as Fig. 3 of higher total confidence on incorrectly answered questions. For each question, we generate a main answer and 10 distractors, so the maximum possible total confidence is $1$.
D.2 $\rm P(True)$} vs. Verbalized Numerical Confidence
@@ D.1 Preliminary Study Extended
In Tables 10 and 11, we expand on Tables 1 and 3 and report results on TriviaQA and biography generation when replacing $\rm P(True)$} with verbalized numerical confidence (see Appendix B.2 for prompts). For SimpleQA, the black-box variants of our methods in Table 2 use verbalized numerical confidence rather than $\rm P(True)$}.
@@
D.3 NLI Model Choice
Table 12: NLI model choice ablation, building on Table 1 which uses the first NLI model. Text styling follows Table 1. DiNCo and SC are robust to the choice of the NLI model, with DiNCo consistently performing the best.
Qwen3-8B Qwen3-1.7B Llama-3.2-3B-Instruct Gemma-3-4B-IT
NLI model: MoritzLaurer/DeBERTa-v3-base-mnli-fever-anli (184M parameters)
NLI model: sileod/deberta-v3-base-tasksource-nli (184M parameters)
SC (Xiong et al., 2024) 0.220 0.241 0.778 0.216 0.224 0.771 0.087 0.195 0.774 0.292 0.300 0.709
NVC 0.114 0.160 0.871 0.084 0.163 0.806 0.155 0.188 0.841 0.157 0.203 0.813
DiNCo 0.108 0.157 0.879 0.189 0.184 0.835 0.052 0.148 0.864 0.137 0.193 0.821
NLI model: MoritzLaurer/mDeBERTa-v3-base-xnli-multilingual-nli-2mil7 (279M parameters)
SC (Xiong et al., 2024) 0.225 0.245 0.771 0.212 0.226 0.765 0.098 0.204 0.757 0.289 0.298 0.712
NVC 0.165 0.190 0.855 0.081 0.173 0.788 0.192 0.203 0.836 0.213 0.229 0.807
DiNCo 0.088 0.159 0.865 0.173 0.182 0.825 0.058 0.152 0.859 0.106 0.190 0.815
NLI model: microsoft/deberta-large-mnli (406M parameters)
SC (Xiong et al., 2024) 0.229 0.242 0.783 0.228 0.225 0.782 0.067 0.180 0.801 0.300 0.303 0.713
NVC 0.141 0.169 0.873 0.085 0.167 0.802 0.168 0.191 0.844 0.191 0.218 0.813
DiNCo 0.103 0.155 0.878 0.178 0.181 0.831 0.045 0.148 0.864 0.121 0.190 0.818
Our main experiments in Section 3 use the NLI model DeBERTa-v3-base-mnli-fever-anli (He et al., 2021). Here we ablate the particular choice of the NLI model. We choose NLI models that are lightweight (0.2-0.4B parameters) and frequently downloaded from HuggingFace. Table 12 confirms that DiNCo and SC are robust to the choice of the NLI model, with DiNCo consistently performing the best.
D.4 Budget Hyperparameters
Figure 8: Calibration at different budget splits (Qwen3-8B on TriviaQA). Setting DiNCo’s (SC sample, distractor) split to the straightforward choice of $(\frac{1}{2},\frac{1}{2})$ as in Section 3 appears to result in the best performance. Nevertheless, DiNCo is robust to the exact budget split, achieving strong calibration across the wide range of budget splits from $(\frac{1}{4},\frac{3}{4})$ to $(\frac{3}{4},\frac{1}{4})$. However, a budget split of $(0,1)$ or $(1,0)$ with only distractors or only SC samples leads to suboptimal calibration, which is not remedied by scaling up the budget due to diminishing returns, empirically supporting our motivation (Section 2.3) for combining the confidence signals from generation and validation.
For our main experiments in Section 3, we gave each method an inference budget of $K=10$, with DiNCo splitting the budget evenly into 5 distractors and 5 SC samples. The equal budget split is a straightforward choice requiring no hyperparameter tuning and achieves strong calibration in Section 3. Here we examine how robust DiNCo is to the budget split and whether tuning it can improve performance. Fig. 8 shows that the budget split can be varied substantially while maintaining performance, revealing DiNCo’s robustness to the budget split hyperparameters. Thus, we recommend the equal budget split for simplicity. However, we find that regardless of the budget, using no SC samples or no distractors performs worse than allocating at least part of the budget to each component, emphasizing the benefit of jointly scaling the inference budgets for generation and validation.
D.5 Open-source LLM Size
Table 13: TriviaQA results with Qwen3-32B. Text styling follows Table 1. The effectiveness of DiNCo extends from the 2-8B scale (Table 1) to the 32B scale.
VC (Kadavath et al., 2022) 0.153 0.158 0.862
$K$-VC (Tian et al., 2023) 0.186 0.211 0.737
MSP (Fadeeva et al., 2023) 0.113 0.164 0.792
SC-VC (Xiong et al., 2024) 0.166 0.213 0.707
SC (Xiong et al., 2024) 0.129 0.190 0.741
NVC 0.194 0.169 0.880
DiNCo 0.065 0.131 0.863
Table 1 shows that DiNCo achieves strong calibration across small to medium-scale (2-8B) open-source models, and Table 2 extends this result to frontier models. Here we consider scaling up within open models, which helps more precisely characterize DiNCo’s performance across scales due to known open model sizes. Table 13 confirms that DiNCo continues to achieve strong calibration at the 32B model size.
D.6 DiNCo in the Medical Domain
We test the generalization of DiNCo to other domains by evaluating on BioASQ (Krithara et al., 2023), a dataset reflecting the information needs of biomedical experts. We sample 1000 factoid questions from the Task B training set of the 2026 challenge edition BioASQ14; see Table 14 for examples. Given the technical expertise required for this task, we evaluate Qwen3-32B, thus also building on Appendix D.5 to confirm DiNCo’s effectiveness on large open-source models. Table 15 shows that the calibration of DiNCo – under the same setup as with TriviaQA – generalizes to the biomedical domain.
Table 14: Example factoid questions and answers from BioASQ.
Question True answer
What is the typical alteration of the thyroid profile metabolism early after coronary artery bypass graft surgery? Low T3 syndrome occurs frequently after CABG
What is the prognostic role of alterred thyroid profile after cardiosurgery? Altered thyroid profile after cardiosurgery is associated with several events in adults and in children
Which is the most common gene signature in Rheumatoid Arthritis patients? Interferon signature
Table 15: BioASQ results with Qwen3-32B. Text styling follows Table 1. DiNCo generalizes to a setting requiring biomedical expertise.
VC (Kadavath et al., 2022) 0.255 0.257 0.801
$K$-VC (Tian et al., 2023) 0.256 0.270 0.749
MSP (Fadeeva et al., 2023) 0.215 0.249 0.743
SC-VC (Xiong et al., 2024) 0.155 0.233 0.710
SC (Xiong et al., 2024) 0.120 0.217 0.727
NVC 0.107 0.167 0.828
DiNCo 0.071 0.160 0.822
D.7 Distractor Generation and Validation in One Step
Our scaling analysis in Section 4 shows that, while DiNCo incurs a slightly higher cost than SC, SC cannot match the calibration of DiNCo even when scaled arbitrarily. Here, we consider further optimizing efficiency by addressing the main reason for DiNCo’s higher cost: the separate step to elicit a verbalized confidence for each distractor. To combine the generation and validation of a distractor into a single conversation, we add a user response containing Output "Yes" if your answer is correct or "No" if not. after the LLM’s answer. Thus, after generating the distractor, only one more input sentence and one more output token need to be processed to obtain a verbalized confidence. We remark, however, that this approach is not applicable to API-access models for which we cannot seamlessly switch between beam search and ordinary decoding, which motivated our two-step design that can be shared between open and closed models. Furthermore, decoupling the generation and validation of distractors allows for other variants, such as using a smaller model to generate distractors for efficiency, as discussed in Appendix C.1. Nonetheless, here we investigate the potential for a one-step design to reduce cost in certain settings while maintaining performance. Table 16 shows that the one-pass variant achieves similar calibration to the two-pass variant, while reducing the relative cost over SC from 32% to 10%.
Table 16: Comparison of DiNCo’s original two-step design (separate calls for distractor generation and validation) and the one-step design (combining both parts into one conversation for efficiency). With Qwen3-8B on 1000 TriviaQA questions, SC uses $9.0\times 10^{15}$ FLOPs, DiNCo (2-step) uses $1.2\times 10^{16}$ (32% more than SC), and DiNCo (1-step) uses $9.9\times 10^{15}$ (10% more than SC). The similar calibration between the two-step and one-step variants of DiNCo suggests that a distractor can be generated and validated jointly to minimize overhead without compromising calibration.
Qwen3-8B Qwen3-1.7B Llama-3.2-3B-Instruct Gemma-3-4B-IT
DiNCo (2-step) 0.097 0.155 0.879 0.177 0.179 0.835 0.044 0.148 0.864 0.121 0.191 0.817
DiNCo (1-step) 0.109 0.171 0.852 0.165 0.191 0.798 0.051 0.157 0.851 0.089 0.210 0.772